# Quasinormal modes in semi-open systems

### v2

@@ Appendix B Pöschl-Teller resonance condition
$\frac{\partial^{2}\psi}{\partial x^{2}}-\left(\omega^{2}-\frac{V_{0}}{\cosh(\alpha x)^{2}}\right)\psi=0\,.$ (18)
@@
$(1-y)y\frac{\partial^{2}\phi}{\partial y^{2}}+[c-(a+b+1)y]\frac{\partial\phi}{\partial y}-ab\phi=0\,,$ (19)
whose solution is (Quasinormal modes in semi-open systems), by introducing the variable:
@@
The parameters $a,b,c$ and the spatial variable $y$ are given in the main text. In order to impose the radiating boundary condition (2) at $x\to-\infty$, one sees that for large negative positions $y^{\pm i\omega/2\alpha}\to e^{\pm i\omega x}$. Hence, it is possible to identify the right-moving contribution in (Quasinormal modes in semi-open systems) and then to impose the boundary condition on the left with $B=0$.

### v3

@@ Appendix A Pöschl-Teller resonance condition
$\frac{\partial^{2}\psi}{\partial x^{2}}-\left(\omega^{2}-\frac{V_{0}}{\cosh(\alpha x)^{2}}\right)\psi=0\,.$ (A.1)
@@
$(1-y)y\frac{\partial^{2}\phi}{\partial y^{2}}+[c-(a+b+1)y]\frac{\partial\phi}{\partial y}-ab\phi=0\,,$ (A.2)
whose solution is (On the origin of quasinormal modes in semi-open systems), by introducing the variable:
@@
The parameters $a,b,c$, and the spatial variable $y$ are given in the main text. In order to impose the radiating boundary condition (2) at $x\to-\infty$, one sees that for large negative positions $y^{\pm i\omega/2\alpha}\to e^{\pm i\omega x}$. Hence, it is possible to identify the right-moving contribution in (On the origin of quasinormal modes in semi-open systems) and then to impose the boundary condition on the left with $B=0$.
@@
Appendix B Numerical simulations
$\partial_{x}^{2}\Psi_{k}=\frac{\Psi_{k+1}-2\Psi_{k}+\Psi_{k-1}}{h^{2}}.$ (B.1)
We then define $\Pi=\partial_{t}\Psi$ and cast (1) in the form,
$\partial_{t}\begin{pmatrix}\Psi\\ \Pi\end{pmatrix}=\begin{pmatrix}0&1\\ \partial_{x}^{2}-V&0\end{pmatrix}\begin{pmatrix}\Psi\\ \Pi\end{pmatrix},$ (B.2)
$\partial_{x}\Psi_{k}=\frac{\Psi_{k+1}-\Psi_{k-1}}{2h},$ (B.3)
$\partial_{x}^{2}\Psi_{1}=\frac{2\Psi_{2}-2\Psi_{1}}{h^{2}}-\frac{2\Pi_{1}}{h}.$ (B.4)
$\partial_{t}\begin{pmatrix}\Psi\\ \Pi\end{pmatrix}=\begin{pmatrix}\mathbf{0}&\mathbf{I}\\ {}^{\mathrm{N}}\partial_{x}^{2}-V&B\end{pmatrix}\begin{pmatrix}\Psi\\ \Pi\end{pmatrix},$ (B.5)
where $B=-(2/h)\mathrm{diag}(1,0,...0,\mu)$ is a diagonal matrix with non-zero entries in the top-left and bottom-right corner. The parameter $\mu$ allows us to tune between a fully reflective Neumann boundary condition on the right grid boundary for $\mu=0$ and a completely open boundary for $\mu=1$. The relation between $\mu$ and the reflectivity $\varepsilon$ used in the main text is,
$\mu=\frac{1-\varepsilon}{1+\varepsilon}.$ (B.6)
To generate noise, we add a source term to the right-hand side of Eq. (B.5) in the form of $(\mathbf{0}\ J(x))^{\mathrm{T}}$. We generate random noise on the open side of the system by setting $J(x)$ everywhere equal to zero except at the point $x_{5}$ where we set $J(x_{5})$ equal to a different random number between $-1$ and $1$ at each time step.